# Explicit Residual-Based Scalable Image Coding for Humans and Machines

### v1

@@ Abstract
Figure 1: Overview of scalable image coding pipelines. (a) Conventional task-specific method, (b) ICMH-FF using decoder-side feature fusion, (c) FR-ICMH introducing encoder-side feature residual modeling, (d) PR-ICMH employing pixel-level residual compression.
@@ II-A Learned Image Compression and Image Coding for Machines
In (1), $y$ and $z$ denote the output of the encoder and hyperprior-encoder of the LIC model, respectively. $\mathcal{R}(y)$ and $\mathcal{R}(z)$ represent the estimated bitrates of $y$ and $z$. $x$ is the original image, and $\hat{x}$ is the reconstructed image. $mse$ denotes the mean squared error function. $\lambda$ is a weighting parameter that controls the trade-off between bitrate and distortion. $m$ is the object mask obtained by the Segment Anything model. The model architecture is based on LIC-TCM with Ch-ARM for entropy modeling.
@@ II-B Scalable Image Coding for Humans and Machines
Scalable coding for human and machine vision has emerged as an important direction in recent research. H. Choi et al. [2] proposed a dual-stream framework that splits the latent representation extracted from the input image into machine- and human-oriented components. These two streams are transmitted separately, and jointly decoded to reconstruct a human-oriented image. More recently, Adapt-ICMH [3] introduced a plug-and-play adaptation module called the spatial-frequency modulation adapter. In this approach, a human-oriented LIC is employed as a base model, and the adapter is inserted and trained to convert it into a machine-oriented codec. Notably, the adapter can be integrated into any LIC regardless of their model structures. Although these methods are effective in terms of scalability and compression performance, they are trained to optimize recognition accuracy and therefore remain task-specific.
@@ II-C Compression Methods Using Residual Connection
Figure 2: Overall architecture of the proposed FR-ICMH, which computes and encodes feature-level residuals between the SA-ICM (upper row) and an additional LIC (lower row). The residuals are subsequently fused with the quantized features from SA-ICM to reconstruct human-oriented images.
@@ III-A Overview
Figure 4: Overall architecture of the proposed PR-ICMH, which computes pixel-level residuals between the machine-oriented output of SA-ICM (upper row) and the original image. The residuals are encoded by an additional LIC (lower row) and added back to reconstruct human-oriented images.
Figure 5: Examples of reconstructed images for machine and human vision. (a) Ground truth, (b) Machine-oriented reconstruction by SA-ICM ($N_{m}=5$), (c) Human-oriented reconstruction by ICMH-FF ($N_{a}=4$), (d) Human-oriented reconstruction by the proposed FR-ICMH ($N_{a}=4$), and (e) Human-oriented reconstruction by the proposed PR-ICMH ($N_{a}=4$). All reconstructions are obtained with $\lambda=0.05$.
@@ III-B Feature Residual-based Scalable Coding
$\hat{y}=conc(\hat{y_{1}},\hat{y_{2}},\ldots,\hat{y_{N_{m}}}).$ (4)
In (3) and (4), $\hat{y}$ denotes the input to the main decoder of additional LIC model. $conc$ stands for the concatenate function. Only the additional LIC model for residual information is trained using the following loss function:
$\mathcal{L}=\mathcal{R}(ya)+\mathcal{R}(z)+\lambda\cdot mse(x,\hat{x}).$ (5)
@@ IV-A Performance of Image Compression for Humans
We evaluate the compression performance for humans of our proposed methods, FR-ICMH and PR-ICMH. For ICM, we utilize SA-ICM model pre-trained with $\lambda=0.05$ according to the loss function defined in (1). For the additional LIC model for residual information, LIC-TCM is utilized. In both methods, only the additional LIC is trained. Since the parameters of SA-ICM are fixed, its compression performance for machines is the same as that of SA-ICM. We train both FR-ICMH and PR-ICMH using their respective loss functions, (5) and (6). During the training, five different $\lambda$ values $\{0.005,0.01,0.02,0.03,0.05\}$ are utilized. COCO-train dataset is utilized for training and COCO-val is utilized for evaluation [42]. The number of feature slices for SA-ICM, $N_{m}$, is fixed to 5, while the additional LIC is trained with $N_{a}=\{3,4,5\}$ to evaluate the effect of number of slices of residual features. Fig. 5 shows that the proposed FR-ICMH and PR-ICMH provide clearer reconstructions of critical visual details, such as facial components and textures in grass, which are less distinct in ICMH-FF.
The rate-distortion curves for human-oriented images with $N_{a}=5$ are presented in Fig. 6. The solid lines indicate the bitrate of additional information only, while dashed lines represent the total bitrate (machine-oriented + additional). We compare our proposed methods with ICMH-FF and LIC-TCM. FR-ICMH outperforms ICMH-FF especially at high bitrate regions, while PR-ICMH consistently outperforms across all bitrate ranges. Notably, PR-ICMH achieves performance that is even close to LIC-TCM at high bitrate levels, despite its scalable structure. The performance degradation of FR-ICMH at low bitrates is primarily due to subtracting the high-quality machine-oriented features of SA-ICM ($\lambda=0.05$) from the human-oriented features. This subtraction yields a very small residual, and results in limited information recovery and lower PSNR.

### v2

@@ Abstract
Figure 1: Overview of scalable image coding pipelines. (a) Conventional task-specific method, (b) ICMH-$F$ using decoder-side feature fusion, (c) FR-ICMH introducing encoder-side feature residual modeling, (d) PR-ICMH employing pixel-level residual compression.
@@ II-A Learned Image Compression and Image Coding for Machines
In (1), $y$ and $z$ denote the output of the encoder and hyperprior-encoder of the LIC model, respectively. $\mathcal{R}(y)$ and $\mathcal{R}(z)$ represent the estimated bitrates of $y$ and $z$. $x$ is the original image, and $\hat{x}$ is the reconstructed image. msemse denotes the mean squared error function. $\lambda$ is a weighting parameter that controls the trade-off between bitrate and distortion. $m$ is the object mask obtained by the Segment Anything model. The model architecture is based on LIC-TCM with Ch-ARM for entropy modeling.
@@ II-B Scalable Image Coding for Humans and Machines
Scalable coding for human and machine vision has emerged as an important direction in recent research. H. Choi et al. [2] proposed a dual-stream framework that splits the latent representation extracted from the input image into machine- and human-oriented components. These two streams are transmitted separately, and jointly decoded to reconstruct a human-oriented image. More recently, Adapt-ICMH [3] introduced a plug-and-play spatial-frequency modulation adapter for converting a human-oriented LIC into a machine-oriented codec. Scalability is achieved by operating either with the adapter for machine-oriented coding or without it for human-oriented coding. Notably, the adapter is compatible with any LIC architecture. Although these methods are effective in terms of scalability and compression performance, they are trained to optimize recognition accuracy and therefore remain task-specific.
@@ II-C Compression Methods Using Residual Connection
Figure 2: Overall architecture of the proposed FR-ICMH, which computes and encodes feature-level residuals between fixed SA-ICM (upper row) and an additional LIC (lower row). The residuals are subsequently fused with the quantized features from SA-ICM to reconstruct human-oriented images.
@@ III-A Overview
Figure 4: Overall architecture of the proposed PR-ICMH, which computes pixel-level residuals between the machine-oriented output of fixed SA-ICM (upper row) and the original image. The residuals are encoded by an additional LIC (lower row) and added back to reconstruct human-oriented images.
Figure 5: Examples of reconstructed images for machine and human vision. (a) Ground truth, (b) Machine-oriented reconstruction by SA-ICM ($N_{m}=5$), (c) Human-oriented reconstruction by ICMH-$F$ ($N_{a}=4$), (d) Human-oriented reconstruction by the proposed FR-ICMH ($N_{a}=4$), and (e) Human-oriented reconstruction by the proposed PR-ICMH ($N_{a}=4$). All reconstructions are obtained with $\lambda=0.05$.
@@ III-B Feature Residual-based Scalable Coding
[…]
In (3) and (4), $\hat{y}$ denotes the input to the main decoder of additional LIC model. concconc stands for the concatenate function. Only the additional LIC model for residual information is trained using the following loss function:
[…]
@@ IV-A Performance of Image Compression for Humans
We evaluate the compression performance for humans of our proposed methods, FR-ICMH and PR-ICMH. For ICM, we utilize fixed SA-ICM model pre-trained with $\lambda=0.05$ according to the loss function defined in (1). For the additional LIC model for residual information, LIC-TCM is utilized. In both methods, only the additional LIC is trained. Since the parameters of SA-ICM are fixed, FR-ICMH and PR-ICMH achieve exactly the same performance as SA-ICM for all recognition tasks, including object detection and segmentation. We train both FR-ICMH and PR-ICMH using their respective loss functions, (5) and (6). During the training, five different $\lambda$ values $\{0.005,0.01,0.02,0.03,0.05\}$ are utilized. COCO-train dataset is utilized for training and COCO-val is utilized for evaluation [42]. The number of feature slices for SA-ICM, $N_{m}$, is fixed to 5, while the additional LIC is trained with $N_{a}=\{3,4,5\}$ to evaluate the effect of number of slices of residual features. Fig. 5 shows that the proposed FR-ICMH and PR-ICMH provide clearer reconstructions of critical visual details, such as facial components and textures in grass, which are less distinct in ICMH-FF.
The rate-distortion curves for human-oriented images with $N_{a}=5$ are presented in Fig. 6. The solid lines indicate the bitrate of additional information only, while dashed lines represent the total bitrate (machine-oriented + additional). We compare our proposed methods with three existing methods: ICMH-FF, LIC-TCM, and VVC (VTM), a traditional compression codec. FR-ICMH outperforms ICMH-FF especially at high bitrate regions, while PR-ICMH consistently outperforms across all bitrate ranges. Notably, PR-ICMH slightly outperforms LIC-TCM at high bitrate levels, despite its scalable structure. The performance degradation of FR-ICMH at low bitrates is primarily due to subtracting the high-quality machine-oriented features of SA-ICM ($\lambda=0.05$) from the human-oriented features. This subtraction yields a very small residual, and results in limited information recovery.